Please Help!!!The slant shear test is widely accepted for evaluating the bond of resinous repair materials to concrete; it utilizes cylinde

Thank you!!!
Answer #1

Here we have: X 1 = 18.10 X2 = 23.17 S1 = 1.59 S2 = 5.02 n1 = 10 n2 = 10 a = 0.05 Hypothesis: Correct option B Ho; M1 – U2

(5 – #z) – (My – uz) = + S2 M ni II = -3.05 (18.10–23.17)-(0) 5.022 + 10 10 1.592 + sz n2 ni df = 2 Si ni $2 n2 + n - 1 n2 -

df 1.592 5.022 + -)2 10 10 1.592 -)2 2 10 + 10 10-1 10-1 = 11 5.022 P-value = 2p (t>/-3.05) = 0.011 Correct option B

Reject Ho. The data suggest that the true average strength is different for the two methods of surface preparation Correct op